Miller Toy Company manufactures a plastic swimming pool at its Westwood Plant. The plant has been experiencing problems as shJanet Dunn, who has just been appointed general manager of the Westwood Plant, has been given instructions to “get things undDuring June, the plant produced 5,000 pools and incurred the following costs: a. Purchased 24,500 pounds of materials at a co1a. Compute the following variances for June, materials price and quantity variances. 1b. Compute the following variances forSummarize the variances that you computed in (1) above by showing the net overall favorable or unfavorable variance for the m
Answer #1

In order to find the reasons for difference between Budgeted and Actual variable cost of good sold we have to compute the material, labour and variable overhead variances.

Solution to 1a

Given That,

Standard Price of Material = $2.5/ pound

Actual Price of Material = $2.95/pound

Standard Quantity of material for 5000 pools = 5,000 * 3.9 pound material/ pool = 19,500 pound

Actual Quantity of material for 5000 pools = 19,300 pound

Material Price Variance = Actual Quantity (Standard Price - Actual Price)

= 19300 (2.5 -2.95)

= (8,685)

Material Quantity Variance = Standard Price (Standard Quantity - Actual Quantity)

= 2.2 (19,500 - 19,300)

= 500

Solution to 1b

Given That,

Actual Direct Lahour Hours Rate = $7.7

Standard Direct Lahour Hours Rate = $8

Actual Direct Labour Hours = 4,600

Standard Direct Labour Hours for 5000 pool = 5,000 * standard direct labour hours for 1 pool

= 5,000 * 0.8

= 4,600

Labour Rate Variance = Actual Hours (Standard Rate - Actual Rate)

= 4600 (8 - 7.7)

=1,380

Labour Efficiency Variance = Standard Rate (standard Hours - Actual Rate)

= 8 (4000 - 4600)

= (4,800)

Solution to 1c

Given That,

Actaual Variable Overhead Machine Hours = 1300

Standard Variable Overhead Machine Hours = 5,000 * Standard Hours for production of 1 pool

= 5,000 * 0.2

= 1,000

Standard Variable Overhead Cost = Standard Variable Overhead Rate * Standard Hours for production of 1 pool * Actual Production

= 3.5*0.2*5,000

= $3,500

Actual Variable Overhead Cost = $5,070

Variable Overhead Rate Variance = Standard cost - Actual Cost

= 3500 - 5070

= (1,570)

Variable Overhead Efficiency Variance = Standard Rate (Standard Hours - Actual Hours)

= 3.5 (1000 - 1300)

= (4,200)

Solution to 2

Material Price Variance 8685 U
Material Quantity Variance 500 F
Labour Rate Variance 1380 F
Labour Efficiency Variance 4800 U
Variable Overhead Rate Variance 1570 U
Variable Overhead Efficiency Variance 4200 U
Net Variance 17,375 U